A 930 kg motorboat accelerates away from a dock at 2.3m?s^2. Its propeller provides a 3.9kN thrust force. What drag force does the water exert on the boat?
1
Expert's answer
2018-05-25T11:41:07-0400
Force equation F=m*a
Force that exert on boat Fb=mb*ab=930*2.3=2139 N
Propeller thrust force Fp is 3.9 kN = 3900 N
Sum all forces taking account its directions
Fb+Fw-Fp=0
Fw=Fp-Fb= 3900 - 2139 = 1761 N
Answer: drag force does the water exert on the boat is 1761 N